OpenStudy (anonymous):

Bees must fly long distances between flowers. Since each bee wants to bring back as much nectar as possible over the course of the day, she needs to maximize the rate of nectar uptake per visit (in microliters/minute), which includes the travel time (in minutes) between flowers. Rate at which nectar is collected = food per visit / total time per visit OR R(Tf) = food per visit [F(Tf)] / [time on flower (Tf) + travel time (Tt)] OR R(Tf) = F(Tf) / (Tf + Tt) where F(Tf) = Tf /( Tf + .5) and Tt = 1 minute. a. Rewrite the equation for R(Tf) using the information given, and simplify. b. Find the firBees must fly long distances between flowers. Since each bee wants to bring back as much nectar as possible over the course of the day, she needs to maximize the rate of nectar uptake per visit (in microliters/minute), which includes the travel time (in minutes) between flowers. Rate at which nectar is collected = food per visit / total time per visit OR R(Tf) = food per visit [F(Tf)] / [time on flower (Tf) + travel time (Tt)] OR R(Tf) = F(Tf) / (Tf + Tt) where F(Tf) = Tf /( Tf + .5) and Tt = 1 minute. a. Rewrite the equation for R(Tf) using the information given, and simplify. b. Find the fir

OpenStudy (nottim):

HI

OpenStudy (nottim):

Finish your question please.

hero (hero):

What the...

hero (hero):

Someone should learn to use Equation editor

OpenStudy (nottim):

What what? LOSASAYWHATINDAWORLDZ

OpenStudy (anonymous):

someone should learn to stop giving smartass remarks

OpenStudy (anonymous):

Bees must fly long distances between flowers. Since each bee wants to bring back as much nectar as possible over the course of the day, she needs to maximize the rate of nectar uptake per visit (in microliters/minute), which includes the travel time (in minutes) between flowers. Rate at which nectar is collected = food per visit / total time per visit OR R(Tf) = food per visit [F(Tf)] / [time on flower (Tf) + travel time (Tt)] OR R(Tf) = F(Tf) / (Tf + Tt) where F(Tf) = Tf /( Tf + .5) and Tt = 1 minute. a. Rewrite the equation for R(Tf) using the information given, and simplify. b. Find the first derivative, R’(Tf). Determine where the original function R(Tf) is increasing and decreasing on the domain [0,3]. Use this information to find the values of any relative extrema.

hero (hero):

You look too pretty to talk smack to me :P

OpenStudy (nottim):

Thx.

OpenStudy (anonymous):

hahah can i please just get help?

hero (hero):

Not Tim should be able to help

OpenStudy (nottim):

I can NOT

OpenStudy (nottim):

attempt to draw the situation with the DRAW button.

hero (hero):

NOtTim, I don't want to hear about what you can NOT do

OpenStudy (anonymous):

if i could i would

OpenStudy (anonymous):

any idea how to graph the function?

OpenStudy (anonymous):

. Graph R(Tf) and R’(Tf) on the domain [0,3]. Don’t forget to label the axes & indicate scale! e. Now how many minutes should the bee spend on each flower? Interpret both graphs in terms of the bee maximizing its time at each flower.

OpenStudy (nottim):

it looks confusing in the format. please attempt to write the equation in a more understandable format.

OpenStudy (anonymous):